OpenStudy (anonymous):

Simplify 10(m + 3) 10m + 3 10m + 30 m + 3 m + 30

OpenStudy (anonymous):

Are those answer choices?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Use the distributive property a(b+c)=ab+ac. Does this help?

OpenStudy (anonymous):

= 10m + 30

OpenStudy (anonymous):

So, you have to remove the brackets and multiply each term by 10. So it will be 10*m+10*3= the second option and also what is written above.